OpenStudy (anonymous):

Which statement is true about the product square root of 2(5 + square root of 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

\[\sqrt2(5+\sqrt8) = 5\sqrt2 + \sqrt8\sqrt2 = 5\sqrt2 + \sqrt{16} = 5\sqrt2 + 4\]

OpenStudy (anonymous):

I used the following identities: \[ a(b+c) = ab + ac\\ \sqrt{a}\sqrt{b} = \sqrt{ab} \]

OpenStudy (anonymous):

but whats the awnser? @JoelSjogren